OpenStudy (anonymous):

A trapezoid is drawn on a grid with vertices at P (-5, 7), Q (-2, 7), R (2, 3), and S (-8, 3). Trapezoid UVWX will be drawn similar to Trapezoid PQRS with side WX corresponding to side RS. The point (8, -4) will be the location of vertex W and point (-12, -4) will be the location of point X. What will be the coordinates of vertex V? Show all your work or explain your reasoning.

OpenStudy (amistre64):

this question is similar to others that have been asked in the past; the key is finding the scalar value between the images i believe

OpenStudy (anonymous):

Ok :) how do i do that?

OpenStudy (amistre64):

since PQ is a straight line; the distance can be determined rather simply

OpenStudy (amistre64):

the distance between -5 and -2 is 3 hmm, but they give you rs and wx, so pq prolly aint much use at the moment :)

OpenStudy (amistre64):

you know how to find the distance between 2 points?

OpenStudy (anonymous):

nope ;(....

OpenStudy (amistre64):

R (2, 3), and S (-8, 3) W (8, -4) and X (-12, -4) subtract one point from the other and pythag the results

OpenStudy (amistre64):

but in this case its a little simpler notice that R and S both end in the same number; which means they differ only in their first numbers

OpenStudy (anonymous):

ok :)

OpenStudy (amistre64):

the distance between -2 and 8 would be: 8--2 = 8+2 = ?

OpenStudy (anonymous):

Can you tell me what the anwser is. so i know what it is and i know how to get to it :)

OpenStudy (amistre64):

the answer will be what we get to in the end; since we are not at the end yet, i wouldnt know what it is :)

OpenStudy (anonymous):

okay:) is it 10 to your question

OpenStudy (amistre64):

10 is the distance from R to S now we find the distance from W to X to compare with since W and X have the same last numbers, the distance is the difference in the first numbers as well: 8 --12 = 8+12 = ?

OpenStudy (anonymous):

20?

OpenStudy (amistre64):

20 is good; now we can find the scaling factor between the traps; 10:20 , or simply put, times 2 since 10*2 = 20

OpenStudy (amistre64):

knowing this, we need to find the distances between the xs and ys of the cooresponding pqrs uvwx v and q coorespond to each other; so the distance from say, r to q is scaled to get w to v

OpenStudy (amistre64):

Q (-2, 7), R (2, 3) we just subtract one point from the other Q (-2, 7) -R (2, 3) -------- (0 , 4) does that make sense?

OpenStudy (anonymous):

is that the anwser?

OpenStudy (amistre64):

not yet, that just tells us how to move from one point to the other; but we are almost there

OpenStudy (amistre64):

this movement (0,4) has to be doubled; so times it by 2 (0,4) * 2 = (0,8) this is the cooresponding movement between W and V just gotta make sure we move in the correct direction

OpenStudy (anonymous):

ok is that the real anwser?

OpenStudy (amistre64):

you really arent following along with this explanation are you .....

OpenStudy (anonymous):

i am im just limited on time.... i have two mins. to finish this assignment or i'll get a zero ;(

OpenStudy (amistre64):

also, i mis mathed the distance from Q to R; shoulda been -2-2 = -4 7-3 = 4 ------------ -4,4 times 2 = -8,8

OpenStudy (amistre64):

W(8, -4) -8 ,8 ------- 0, 4 W(8, -4) 8,-8 ------ 16,-12 depending on which direction we need to go, its one of these

OpenStudy (amistre64):

you got answer options to compare with?

OpenStudy (anonymous):

no i don't :(

OpenStudy (amistre64):

P (-5, 7), Q (-2, 7), R (2, 3), and S (-8, 3)